ANALYSIS OF LOW FREQUENCY VIBRATIONAL PROGRESSIONS IN THE HIGH RESOLUTION ELEC-TRONIC SPECTRUM of 4,4’-DIMETHYLAMINOBENZONITRILE (DMABN).^{a}

\underline{ALEXEI E.NIKOLAEV}, JAMES F.PFANSTIEL, DAVID W.PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, Pennsylvania 15260, USA; GRZEGORZ MYSZKIEWICZ, GIEL BERDEN^{b}, W. LEO MEERTS, Department of Molecular and Laser Physics, NSRIM, University of Nijmegen, P.O. Box 9010, 6500 GL Nijmegen, The Netherlands.

Recent developments in the study of the DMABN in the gas phase^{c}^{,}^{d} prompted us to reexamine its electronic spectrum at full rotational resolution. This examination shows that several members of the low frequency vibronic progression at +76, +113, +118, +136, +176, +190, and +193 cm^{−1}are split into two or more components. This splitting, not reported before, is due to the varying magnitude of coupling between two internal -CH_{3} rotors in the DMABN molecule in its different vibrational states. The assignment of the observed spectra was greatly facilitated by the use of genetic algorithm automated assignments.^{e}^{,}^{f} Examination of the obtained rotational parameters leads to the determination of the geometry of the molecule in its ground and electronically excited states. There exists a controversy in the literature on the lifetime of the excited electronic state of DMABN between various experiments. The picosecond study^{g} obtained values between 1.5 ps and 24 ps, while the other works^{h} yielded ∼5 ns. From the observed Lorentzian component of the linewidth we determined the lifetime of the excited state to be 4 ± 0.5 ns.

aWork supported by NSF

bPresent address: FOM Institute for Plasma Physics Rijnhuizen, Nieuwegein, The Netherlands

cG. Myszkiewicz, G. Berden, W. Leo Meerts, paper TD01 at 58^{th} OSU International symposium on molecular spectroscopy

dH. Saigusa, N. Miyakoshi, Ch. Mukai, T. Fukagawa, Sh. Kohtani, R. Nakagaki, and R. Gordon J.Chem.Phys 119, 5414 (2003)

eJ.A. Hageman, R. Wehrens, R. de Gelder, W.L. Meerts and L.M.C. Buydens. J. Chem. Phys. 113, 7955 (2000)

fW. L. Meerts, M. Schmitt and G. C. Groenenboom. accepted Can. J. Chem. (2004)

gF. P. Salgado, J. Herbich, A. G. M. Kunst, and R. P. H. Rettschnick, J. Phys. Chem. A 103, 3184 (1999)

hB. R. Howells, J. McCombie, T. F. Palmer, J. P. Simons, and A. Walters J. Chem. Soc., Faraday Trans. 88, 2595 (1992); E.

M. Gibson, A. C. Jones, and D. Phillips Chem. Phys. Lett. 136, 454 (1987); E. M. Gibson, A. C. Jones, D. Phillips, A. G. Taylor, W. G. Bouwman, and D. Sandell, J. Phys. Chem. 92, 5449 (1988)